OpenStudy (anonymous):

Find the coordinates for the midpoint of the segment with endpoints given. (10, 6) and (-4, 8) (7, 7) (3, 7)

hero (hero):

The midpoint formula is \[M(x,y) = \left(\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}\right)\]

OpenStudy (anonymous):

o the answer is number 2 right ? thats what im getting

hero (hero):

The solution will be a point (x,y) of course.

OpenStudy (anonymous):

so im geussing number 1

hero (hero):

@gabby122, the solution will be a point in the form (x,y). Saying the "the answer is 2" or "the answer is 1" is not the solution.

hero (hero):

If you are going to state what the solution is, you should state it explicitly

hero (hero):

For example, if I had to find the midpoint of (2,4) and (4,8) I would state that the midpoint is (3,6). Or that the solution is the point (3,6).

hero (hero):

If I stated the solution is 2, then it wouldn't make sense. The solution must be a point in the form (x,y).